OpenStudy (anonymous):

Find the equations of the tangent plane and normal line to the surface x + y + z = e^(xyz) at the point (0,0,1).

OpenStudy (anonymous):

I think this is what should be done: \[F_x (x,y,z)= 1-yze^{xyz}\] \[F_x (0,0,1)= 1\] \[F_y (x,y,z)= 1-xze^{xyz}\] \[F_y (0,0,1)= 1\] \[F_z (x,y,z)= 1-xye^{xyz}\] \[F_z (0,0,1)= 1\] Then the tangent plane is: \[1(x-0)+1(y-0)+1(z-1)=0\] \[x+y+z=1\] The normal line is: \[x=y=z-1\]